### v3

# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

Abstract.
Motivated by the work of Cheng–Fang–Wang–Yu on the hypersingular Bergman projection, we develop a real-variable and dyadic framework for hypersingular operators in regimes where strong-type estimates fail at the critical line. The main new input is a hypersingular sparse domination principle combined with Bourgain’s interpolation method, which provides a flexible mechanism to establish critical-line (and endpoint) estimates.
In the unit disc setting with $1<t<3/2$, we obtain a full characterization of the $(p,q)$ mapping theory for the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$, in particular including estimates on the critical line 1/q1/p=2t21/q-1/p=2t-2 and a weighted endpoint criterion in the radial setting. We also prove endpoint estimates for the hypersingular Bergman projection
$K_{2t}f(z)=\int_{\mathbb{D}}\frac{f(w)}{(1-z\overline{w})^{2t}}\,dA(w),$
including a restricted weak-type bound at $(p,q)=\bigl(\tfrac{1}{3-2t},1\bigr)$. Finally, we introduce a class of hypersingular cousin of sparse operators in $\mathbb{R}^{n}$ associated with graded sparse families, quantified by the sparseness $\eta$ and a new structural parameter (the degree) $K_{\mathcal{S}}$, and we characterize the corresponding strong/weak/restricted weak-type regimes in terms of $(n,t,\eta,K_{\mathcal{S}})$.
Our real-variable perspective addresses to an inquiry raised by Cheng–Fang–Wang–Yu on developing effective real-analytic tools in the hypersingular regime for $K_{2t}$, and it also provides a new route toward the critical-line analysis of Forelli–Rudin type operators and related hypersingular operators in both real and complex settings.
Key words and phrases:
Hypersingular maximal operator, Hypersingular Bergman projection, Hypersingular sparse operator, graded family, critical line estimates
2020 Mathematics Subject Classification:
42B25, 42B20, 30H20, 46B70, 42B35
1. Introduction
The present paper is motivated by the recent work of Cheng, Fang, Wang, and Yu [12], who studied the following Bergman-type operator on the unit disc $\mathbb{D}$}: for $t>0$,
(1.1) $K_{2t}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}\,dA(w),$
where $dA$ denotes the normalized area measure on $\mathbb{D}$}. The study of the operator $K_{2t}$ can, in broad terms, be divided into three regimes:
  1. (1)
    $t=1$, in which case $K_{2t}=K_{2}$ coincides with the Bergman projection on $\mathbb{D}$};
  2. (2)
    $0<t<1$, in which case $K_{2t}$ is the fractional Bergman projection on $\mathbb{D}$};
  3. (3)
    $t>1$, in which case $K_{2t}$ becomes the hypersingular Bergman projection on $\mathbb{D}$}.
In the first two regimes, the situation is fairly well understood. One key reason is that the Bergman projection can be viewed as a generalized Calderón–Zygmund operator (see, e.g., [20]). Consequently, one may bring to powerful tools from Calderón–Zygmund theory and dyadic harmonic analysis (such as sparse domination) in the study of the Bergman projection, as well as its fractional counterparts. This has proved to be a fruitful line of research, encompassing (among many other directions)
  1. $\bullet$
    weight theory [5, 24, 25, 26],
  2. $\bullet$
    Bergman theory in several complex variables [4, 10, 17, 21, 23, 27], and
  3. $\bullet$
    commutator and BMOBMO theory [9, 15, 19, 30].
We emphasize that the literature in each of these directions is extensive, and the above list is far from exhaustive and is included only for the reader’s convenience.
For the third regime $t>1$, to the best of our knowledge, the existing results rely mainly on complex and functional analytic methods, more precisely within the framework of Forelli–Rudin type operators, dating back to the early work of Forelli and Rudin [11] in 1974. We refer the reader to [12, 28] for more recent developments.
A common feature of these results is that they provide strong $L^{p}$$L^{q}$ bounds only away from the critical line, while estimates on the critical line itself appear to be unavailable in the literature. One main reason is that strong-type bounds on the critical line generally fail (see, Figure 2). This obstruction, in turn, make it difficult to use the techniques that are effective in the case $t\leq 1$, since $K_{2t}$ becomes more singular in the hypersingular regime (see the discussion after (1.3)). Moreover, for general hypersingular operators, weak-type bounds may even fail at one of the endpoints on the critical line (see Remark 1.7), indicating that genuinely new ideas are required to treat critical-line estimates.
Remark 1.1.
It is important to distinguish the hypersingular operators studied in this paper from the strongly singular Calderón–Zygmund operators [^1] The authors thank Árpád Bényi for pointing out this distinction. introduced by Alvarez and Milman [2]. The latter are formulated under the a priori assumption that the operator extends boundedly on $L^{2}(\mathbb{R}^{n})$, together with suitable off-diagonal regularity of the kernel.
In contrast, the hypersingular operators considered here do not admit an $L^{2}$ theory; moreover, strong-type bounds typically fail not only at the diagonal endpoint but even along the critical line (see, e.g., Figures 1, 2, and 3 below). Consequently, hypersingular operators exhibit a genuinely different type of singular behavior from the strongly singular Calderón–Zygmund class.
Remark 1.2.
In [12, Section 1], the authors remarked that, in the hypersingular regime $t>1$, developing a workable real-variable approach that yields explicit boundedness results for concrete operators remains an open direction. This is one of the main motivations for the present work.
The first goal of the current paper is to develop harmonic-analytic methods for the study of the hypersingular Bergman projections $K_{2t}$ for $t>1$, thanks to the recent development in dyadic harmonic analysis. In particular, we aim to establish estimates on the critical line, which, to the best of our knowledge, are even new in the unit disc setting. In this way, we address an inquiry raised in [12] by developing a real–variable approach that yields explicit bounds in the hypersingular regime.
Let us now turn to some details. Instead of studying $K_{2t}$ for $t>1$ directly, it is natural from the viewpoint of harmonic analysis to first consider its maximal analogue, which we call the hypersingular maximal operator.
Definition 1.3.
For $t>1$, the hypersingular maximal operator $\mathcal{M}_{t}$ is defined by
$\mathcal{M}_{t}f(z):=\sup_{\begin{subarray}{c}I\subseteq\mathbb{T}\\ z\in Q_{I}\end{subarray}}\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|\,dA(w),$
where $\mathbb{T}$ denotes the unit circle and $Q_{I}$ is the Carleson box associated with an arc $I\subseteq\mathbb{T}$. It is also convenient to introduce its dyadic analogue. Let $\mathcal{D}$ be a dyadic system on $\mathbb{T}$. For $t>1$, the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$ is defined by
$\mathcal{M}_{t}^{\mathcal{D}}f(z):=\sup_{\begin{subarray}{c}I\in\mathcal{D}\\ z\in Q_{I}\end{subarray}}\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|\,dA(w).$
Remark 1.4.
In what follows, we will restrict our attention to the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$. It is a standard fact (via the $1/3$-trick) that $\mathcal{M}_{t}$ is pointwise comparable to the sum of two dyadic counterparts: there exist two dyadic systems $\mathcal{D}$ and $\widetilde{\mathcal{D}}$ on $\mathbb{T}$ such that
$\mathcal{M}_{t}f(z)\simeq\mathcal{M}_{t}^{\mathcal{D}}f(z)+\mathcal{M}_{t}^{\widetilde{\mathcal{D}}}f(z),\quad z\in{\mathbb{D}}.$
Here $\mathcal{D}$ may be taken to be the standard dyadic system on $\mathbb{T}$, and $\widetilde{\mathcal{D}}$ a $1/3$-shifted dyadic system. This reduction goes back to Mei’s work on $BMO$ and its dyadic analog [22].
We first observe that it only makes sense to consider the case when $1<t<3/2$ in the hypersingular regime. Indeed, it is easy to check that
$\left({\mathcal{M}}_{t}^{{\mathcal{D}}}1\right)(z)=\sup_{z\in Q_{I},\;I\in{\mathcal{D}}}\frac{|Q_{I}|}{|Q_{I}|^{t}}\simeq\frac{1}{(1-|z|^{2})^{2(t-1)}},\qquad z\in{\mathbb{D}}.$
The minimal requirement here is to make ${\mathcal{M}}_{t}^{{\mathcal{D}}}1\in L^{1}({\mathbb{D}})$, and hence
(1.2) $\int_{{\mathbb{D}}}\left|\left({\mathcal{M}}_{t}^{{\mathcal{D}}}1\right)(z)\right|dA(z)\simeq\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})^{2(t-1)}}dA(z)\simeq\int_{0}^{1}\frac{dr}{(1-r)^{2(t-1)}},$
which is finite if and only if $t<3/2$.
Remark 1.5.
The behavior of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ is more subtle than that of the classical maximal operator over Carleson tents and its fractional analogue. In particular, it depends not only on the underlying dyadic structure, but also on the geometry of the ambient domain. More precisely:
  1. (1)
    In the real-variable setting (namely, for the classical Hardy–Littlewood maximal operator), it is clear that it is not meaningful to consider $t>1$, since in that range the operator is not even well-defined on nonzero constant functions.
  2. (2)
    Likewise, if one replaces $\mathbb{D}$} by the upper half plane $\mathcal{H}:=\{z\in\mathbb{C}:\textnormal{Im}z>0\}$, then a careful inspection of the argument in (1.2) yields that the operator is again not well-defined on nonzero constant functions whenever $t>1$ (see also [12, Theorem 5] for a related statement concerning the hypersingular Bergman projection $K^{\mathcal{H}}_{2t}$ on the upper half plane).
We have the following full characterization for the behavior of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$, including the critical line estimates. Let $1<t<3/2$ and $\mathcal{D}$} be any dyadic system on $\mathbb{T}$. Then ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ (as well as ${\mathcal{M}}_{t}$) is
  1. (1)
    (Strong-type bound) bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ for
    $\frac{1}{q}-\frac{1}{p}>2t-2,\quad\textrm{with}\ 1\leq p,q\leq+\infty$
    (see, Proposition 3.3);
  2. (2)
    (Weak-type bound) bounded from $L^{p}({\mathbb{D}})$ to $L^{q,\infty}({\mathbb{D}})$ for
    $\frac{1}{q}-\frac{1}{p}=2t-2,\quad\textrm{with}\ 1\leq p,q\leq+\infty$
    (see, Lemma 3.4, Theorem 3.5, and Corollary 3.7).
The results above are summarized in Figure 1.
$\frac{1}{p}11$$\frac{1}{q}$(0, 2t2)(0,\,2t-2)(32t, 1)(3-2t,\,1)
Figure 1. Boundedness of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ for $1<t<3/2$: the red line and the shaded region indicate strong $(p,q)$ bounds, while the blue line indicates weak $(p,q)$ bounds.
Moreover, at the endpoint $\left(p,q\right)=\left(\frac{1}{3-2t},1\right)$, we obtain necessary and sufficient conditions for the endpoint bounds of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ in the radial weighted setting $\omega(z)=\omega(|z|)$, $z\in{\mathbb{D}}$. More precisely, we prove the following characterizations:
  1. (1)
    Weak-type bounds characterization (see, Theorem 3.5):
    ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}},\omega)\to L^{q,\infty}({\mathbb{D}})\ \textrm{is bounded}\ \Longleftrightarrow\ \sup_{k\geq 0}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\omega(r)^{-\frac{3-2t}{2t-2}}\,dr<+\infty$
  2. (2)
    Strong-type bounds characterization (see, Theorem 3.8): assume in addition that $\omega\in{\bf B}_{\frac{1}{3-2t}}$. Then
    ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}},\omega)\to L^{q}({\mathbb{D}})\ \textrm{is bounded}\ \Longleftrightarrow\ \sum_{k\geq 0}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\omega(r)^{-\frac{3-2t}{2t-2}}\,dr<+\infty.$
    Here, ${\bf B}_{\frac{1}{3-2t}}$ denotes the Békollé–Bonami weight class (see, (3.11)).
Our proofs of the above results build on recent developments in dyadic harmonic analysis, together with underlying geometry properties of dyadic Carleson boxes.
Remark 1.6.
As pointed out in [12] (see the discussion under “Maximal operators” there), the maximal operator ${\mathcal{M}}_{t}$ (or ${\mathcal{M}}_{t}^{{\mathcal{D}}}$) associated with Carleson boxes on $\mathbb{D}$} arises naturally as a hypersingular analogue of the classical maximal operator, and one expects a corresponding boundedness theory in the range $1<t<3/2$. The results above therefore provide such a theory from a real-variable perspective.
Next, we consider the behavior of $K_{2t}$. As observed in [12], it suffices to restrict to the range $1<t<3/2$. In this regime, we obtain a full characterization of the behavior of $K_{2t}$, including the critical-line behavior. In particular, $K_{2t}$ is
  1. (1)
    (Strong-type bound) bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ for
    $\frac{1}{q}-\frac{1}{p}>2t-2,\quad\textrm{with}\ 1\leq p,q\leq+\infty$
    (see, [12, Theorem 3] and also Theorem 5.7, (1) for a generalization and strengthening of this result in terms of “hypersingular” sparse operators);
  2. (2)
    (Weak-type bound) bounded from $L^{p}({\mathbb{D}})$ to $L^{q,\infty}({\mathbb{D}})$ for
[…]
    (see, Lemma 4.1 and Corollary 4.5).
  3. (3)
    (Restricted weak-type bound) bounded from $L^{p,1}({\mathbb{D}})$ to $L^{q,\infty}({\mathbb{D}})$ for
    (1.3) $p=\frac{1}{3-2t}\quad\textrm{and}\quad q=1.$
    (see, Theorem 4.4).
We summarize the above results for $K_{2t}$ in Figure 2.
$\frac{1}{p}11$$\frac{1}{q}$(0, 2t2)(0,\,2t-2)(32t, 1)(3-2t,\,1)
Figure 2. Boundedness of $K_{2t}$ for $1<t<3/2$: the red line and the shaded region indicate strong $(p,q)$ bounds, the blue line indicates weak $(p,q)$ bounds, and the green dot indicates restricted $(p,q)$ bounds.
One major difficulty in analyzing $K_{2t}$ arises at the endpoint $p=\frac{1}{3-2t}$ and $q=1$. Indeed, as observed in [12, Theorem 3], strong-type estimates fail on the critical line 1/q1/p=2t21/q-1/p=2t-2. This is in sharp contrast with the Bergman projection $K_{2}$, whose $L^{2}$ boundedness is simply guaranteed by its definition. As a consequence, both the Calderón–Zygmund approach (for instance, the Calderón–Zygmund decomposition, which uses an $L^{r}$ bound, $1<r<\infty$, to control the bad function) and more recent sparse domination techniques (e.g. the weak-type machinery in [8, Theorem E]) become more delicate to implement directly in this setting.
To address this issue, we develop a new approach based on Bourgain’s interpolation trick (see, Lemma 4.3).
Key idea. Observe that the point $\left(1/p,1/q\right)=(3-2t,1)$ lies between the two corner points $(0,1)$ and $(1,1)$, which correspond to a “good” strong-type regime and a “bad” blow-up regime, respectively. Our strategy is to connect these two behaviors by combining a dyadic/sparse domination perspective with Bourgain’s interpolation trick. Remarkably, this interpolation lands exactly at $\left(1/p,1/q\right)=(3-2t,1)$ and, together with off-diagonal Marcinkiewicz interpolation, yields weak-type bounds along the critical line away from this endpoint.
Remark 1.7.
  1. (1)
    A careful examination of the proof of the above results for $K_{2t}$ shows that the same bounds also hold for the associated positive operator
    (1.4) $K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w),$
    which can be regarded as a hypersingular analogue of the Berezin transform.
  2. (2)
    One may ask whether, at the endpoint (1/p,1/q)=(32t,1)(1/p,1/q)=(3-2t,1), the restricted weak-type estimate (1.3) can be upgraded to a weak-type bound (recall that $\mathcal{M}_{t}^{\mathcal{D}}$ does satisfy a weak-type estimate at this endpoint). The situation, however, is more subtle. Indeed, one can show that the hypersingular Berezin transform (1.4) is unbounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$ (see, Proposition 4.2). For $K_{2t}$, we conjecture that the weak-type estimate also fails at this endpoint (see Section 6 for further discussion).
Motivated by our approach the critical-line estimate for $K_{2t}$, it is natural to ask whether Bourgain’s interpolation method can be used to study hypersingular analogues of sparse operators in harmonic analysis. This leads to the second main goal of the present paper. More precisely, let $t>1$ and let $\mathcal{S}$} be a sparse family in $\mathbb{R}^{n}$ such that [^2] This global containment assumption is natural in the hypersingular setting (see Remark 5.3 for further discussion). there exists a dyadic cube $Q_{0}$ with $Q\subseteq Q_{0}$ for all $Q\in{\mathcal{S}}$. Consider the following hypersingular averaging operator
$\mathbb{A}_{{\mathcal{S}}}^{t}f(x):=\sum_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t}}\int_{Q}|f(y)|\,dy.$
We are interested in determining the admissible $(p,q)$-range for which $\mathbb{A}_{{\mathcal{S}}}^{t}$ is of strong type, weak type, or restricted weak type. Note that the study of $\mathbb{A}_{{\mathcal{S}}}^{t}$ is also of independent interest from the viewpoint of dyadic harmonic analysis, since it can be viewed as a hypersingular counterpart of the classical sparse operator.
It turns out that the boundedness behavior of $\mathbb{A}_{{\mathcal{S}}}^{t}$ is determined by the following four parameters:
  1. (1)
    $n\geq 1$, the real dimension of the ambient space;
  2. (2)
    $t>1$, the hypersingular index of the averaging operator $\mathbb{A}_{{\mathcal{S}}}^{t}$;
  3. (3)
    $\eta\in(0,1)$, the sparseness of $\mathcal{S}$}, measuring how much of each cube can be chosen disjointly (see, Definition 5.1).
  4. (4)
    $K_{{\mathcal{S}}}\geq 1$, the degree of $\mathcal{S}$}, rouphly speaking, which measures the maximal dyadic scale drop between consecutive layers (see, Definition 5.5).
We note that, in the hypersingular regime, it is pivotal to assume that the degree $K_{{\mathcal{S}}}$ is finite. We have the following motivating example.
Example 1.8.
Let $t>1$ and
${\mathcal{S}}_{m}:=\left\{I_{k}:=\left[\frac{k}{2^{m}},\frac{k+1}{2^{m}}\right),0\leq k\leq 2^{m}-1\right\}\cup\left\{[0,2)\right\}.$
for each $m\geq 1$. It is clear that ${\mathcal{S}}_{m}$ is an $\eta$-sparse family for any $\eta\in(1/2,1)$ and any $m\geq 1$; however
$\mathbb{A}_{{\mathcal{S}}_{m}}^{t}1(x)\geq\sum_{k=0}^{2^{m}-1}\frac{\mbox{1\hskip-4.25pt{1}}_{I_{k}}(x)}{|I_{k}|^{t}}\int_{I_{k}}1dx=2^{m(t-1)},\quad x\in[0,1),$
which implies $\left\|\mathbb{A}_{{\mathcal{S}}_{m}}^{t}\right\|_{L^{\infty}([0.2))\to L^{1}([0,2))}\gtrsim 2^{m(t-1)}$. This example shows that, in order to capture the behavior of $\mathbb{A}_{{\mathcal{S}}}^{t}$, the sparseness of $\mathcal{S}$} alone does not suffice, and one must have certain control for the dyadic scales between consectutive layers in $\mathcal{S}$}. This simple example also highlights a fundamental difference between the sparse operator and its hypersingular counterpart: in the former case, the degree plays essentially no role. We refer the reader to Section 5.1 for further discussion of this topic.
Here is the main result in the second part of the paper. Let $\mathcal{S}$} be a sparse family in $\mathbb{R}^{n}$ as above, with sparseness $\eta\in(0,1)$ and degree $K_{{\mathcal{S}}}\in[1,\infty)$; we refer to such a family as a graded sparse family. Then for any $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$, we have $\mathbb{A}_{{\mathcal{S}}}^{t}$ is
  1. (1)
    (Strong-type bound) bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ for
    $\frac{1}{q}-\frac{1}{p}>\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)},\quad\textrm{with}\ 1\leq p,q\leq+\infty;$
  2. (2)
    (Weak-type bound) bounded from $L^{p}({\mathbb{D}})$ to $L^{q,\infty}({\mathbb{D}})$ for
[…]
  3. (3)
    (Restricted weak-type bound) bounded from $L^{p,1}({\mathbb{D}})$ to $L^{q,\infty}({\mathbb{D}})$ for
    $p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\quad\textrm{and}\quad q=1.$
    However, in general the endpoint weak-type estimate for $\mathbb{A}_{{\mathcal{S}}}^{t}$ fails. Indeed, this follows from Proposition 4.2 together with the fact that $K_{2t}^{+}$ can be bounded above pointwisely by a sum of two hypersingular sparse operators (see (4.1)).
We will prove these estimates in Theorem 5.7. For the reader’s convenience, we summarize the resulting bounds below (see Figure 3).
$\frac{1}{p}11$$\frac{1}{q}$$\left(0,\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}\right)$$\left(\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},1\right)$
Figure 3. Boundedness of $\mathbb{A}_{{\mathcal{S}}}^{t}$ for $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$: the red line and the shaded region indicate strong $(p,q)$ bounds, the blue line indicates weak $(p,q)$ bounds, and the green dot indicates restricted $(p,q)$ bounds.
Remark 1.9.
  1. (1)
    A noteworthy feature of the main theorems above is that the geometric parameters of the family—in particular, the sparseness $\eta$ and the degree $K_{{\mathcal{S}}}$—enter the admissible $(p,q)$-range through the critical relation
    $\frac{1}{q}-\frac{1}{p}=\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}.$
    In other words, the geometry of the underlying collection directly affects the range of $L^{p}$$L^{q}$ boundedness for $\mathbb{A}_{{\mathcal{S}}}^{t}$. This is in sharp contrast with the classical sparse operators (corresponding to $t=1$), where sparseness influences only the operator norm, while the set of admissible exponents $(p,q)$ is independent of the particular sparse family.
  2. (2)
    The results above strengthen our earlier characterization of the hypersingular Bergman projection $K_{2t}$ in the range $1<t<3/2$. Indeed, the collection of all dyadic Carleson boxes on $\mathbb{D}$} (in real dimension $n=2$) forms a graded sparse family with sparseness $\eta=1/2$ and degree $K_{{\mathcal{S}}}=1$. With these parameters, one has
    $\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}=2t-2,\quad\text{and}\quad\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}=\frac{1}{3-2t},$
    which exactly coincide with those arising in the mapping theory of $K_{2t}$.
  3. (3)
    The strong-type range above is sharp, as can be seen from the strong-type characterization for $K_{2t}$ (see [12, Theorem 3]). This suggests that the $L^{p}$$L^{q}$ mapping properties of $K_{2t}$ are governed primarily by the underlying geometry of the (dyadic) Carleson boxes, rather than by analyticity.
  4. (4)
    In the results above, the sparseness parameter $\eta$ is understood with respect to the underlying dyadic grid. This is the reason that the quantity $\log_{2}(1-\eta)$ (and hence a base-$2$ logarithm) appears naturally in our main theorem. For a discussion of how this normalization behaves under changing the base of the grid (e.g. dyadic versus triadic systems), see Remark 5.9.
We conclude the introduction by summarizing the main novelties of the present paper, which can be divided into the following aspects.
  1. (a)
    From the viewpoint of complex function theory, we advance the study of Forelli–Rudin type operators in the hypersingular regime by establishing new endpoint and critical-line estimates for $K_{2t}$ with $t>1$. To the best of our knowledge, these estimates are new even in the unit disc setting. A key feature of our approach is the introduction of Bourgain’s interpolation method into this setting, which allows us to overcome the lack of strong-type bounds and provides a robust harmonic-analytic framework for hypersingular operators arising in complex variables.
  2. (b)
    From the perspective of harmonic analysis, we introduce a new dyadic model, which we refer to as a hypersingular sparse operator. Unlike classical sparse operators associated with Calderón–Zygmund theory or recent sparse domination frameworks, the $(p,q)$-boundedness range of these hypersingular sparse operators depends intrinsically on the geometry of the underlying sparse family. In particular, geometric parameters such as sparseness and degree directly influence the admissible $(p,q)$-range, rather than merely affecting operator norms. This reveals a genuinely new phenomenon beyond both classical Calderón–Zygmund theory and standard sparse domination.
  3. (c)
    More broadly, the Bourgain interpolation framework developed in this paper provides a flexible and systematic tool for treating hypersingular operators. We expect that this approach will have further applications, including weighted estimates, commutator theory, hypersingular Bergman theory in several complex variables (on general domains), and related problems, both in complex analysis and in other areas of analysis; see Section 6 for possible directions for future work.
The rest of this paper is organized as follows. In Section 2, we collect basic definitions and notation that will be used throughout the paper. In Section 3, we establish the $L^{p}$ mapping theory for the dyadic hypersingular maximal operator $M_{t}^{\mathcal{D}}$, including both the off-critical-line bounds and the critical-line estimates, together with the endpoint characterizations in the radial weighted setting. In Section 4, we turn to the hypersingular Bergman projection $K_{2t}$ and prove the critical-line estimates, with special emphasis on the endpoint $(p,q)=\left(\frac{1}{3-2t},1\right)$ via a combination of hypersingular sparse domination and Bourgain’s interpolation trick. In Section 5, we introduce hypersingular sparse operators associated with graded sparse families in $\mathbb{R}^{n}$ and prove the corresponding $L^{p}$-$L^{q}$ theory, which in particular extends and strengthens the results obtained earlier for $K_{2t}$. Finally, in Section 6, we propose several open problems and further directions motivated by the methods developed in this paper.
Throughout this paper, for $\in\mathbb{R}, aba\lesssim b$ means there exists a positive number $C$, which is independent of $a$ and $b$, such that $a\leq C\,b$. Moreover, if both $a\lesssim b$ and $b\lesssim a$ hold, then we say $a\simeq b$.
Acknowledgement. The authors are grateful to Árpád Bényi for drawing their attention to strongly singular Calderón–Zygmund operators. They also thank Yongsheng Han, Cody Stockdale, Yun–Hao Lee, Kenan Zhang, and Zipeng Wang for helpful discussions during various stages of this work. The first author was supported by the Simons Travel grant MPS-TSM-00007213.
2. Notations
In this section, we collect several basic definitions and notations that will be used frequently throughout the paper.
Let $I\subseteq\mathbb{T}$ be an arc. Define the Carleson box associated with $I$ by
$Q_{I}:=\left\{z\in{\mathbb{D}}:\ \frac{z}{|z|}\in I,\ 1-|I|\leq|z|<1\right\},$
and also the corresponding upper-half tent by
$Q_{I}^{\mathrm{up}}:=\left\{z\in{\mathbb{D}}:\ \frac{z}{|z|}\in I,\ 1-|I|\leq|z|<1-\frac{|I|}{2}\right\}.$
A dyadic system (or dyadic grid) $\mathcal{D}$} on $\mathbb{T}$ is a collection of arcs
${\mathcal{D}}=\bigcup_{k\geq 0}{\mathcal{D}}_{k},$
where each generation ${\mathcal{D}}_{k}$ consists of $2^{k}$ disjoint arcs of equal length $2^{-k}$ whose union is $\mathbb{T}$ (equipped with the normalized arc measure) and such that each arc $I\in{\mathcal{D}}_{k}$ is the union of two arcs $I^{(1)},I^{(2)}\in{\mathcal{D}}_{k+1}$ (called the dyadic children of $I$). Equivalently, after the identification $\mathbb{T}\simeq[0,1)$, one may take
(2.1) ${\mathcal{D}}_{k}=\Bigl\{\Bigl[\frac{m}{2^{k}},\frac{m+1}{2^{k}}\Bigr):\ m=0,1,\dots,2^{k}-1\Bigr\},$
and view each such interval as an arc on $\mathbb{T}$.
Next, we record a few basic estimates and conventions. Let $1\leq p,q\leq+\infty$, and let $\mathcal{T}$ denote a sublinear operator on $\mathbb{D}$}. The strong-type bound $\mathcal{T}:L^{p}({\mathbb{D}})\to L^{q}({\mathbb{D}})$ is understood in the usual sense.
For $1\leq p<\infty$, the weak Lebesgue space $L^{p,\infty}({\mathbb{D}})$ consists of all measurable functions $f$ on $\mathbb{D}$} such that
$\|f\|_{L^{p,\infty}({\mathbb{D}})}:=\sup_{\lambda>0}\,\lambda\,\big|\{z\in{\mathbb{D}}:\ |f(z)|>\lambda\}\big|^{1/p}<\infty.$
The Lorentz space $L^{p,1}({\mathbb{D}})$ is defined by
$\|f\|_{L^{p,1}({\mathbb{D}})}:=p\int_{0}^{\infty}\lambda^{p-1}\big|\{z\in{\mathbb{D}}:\ |f(z)|>\lambda\}\big|\,d\lambda<\infty,$
with the usual modification when $p=\infty$.
We say that $\mathcal{T}$ is of weak type $(p,q)$, and write $\mathcal{T}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded, if there exists $C>0$ such that for all $f\in L^{p}({\mathbb{D}})$ and all $\lambda>0$,
$\big|\{z\in{\mathbb{D}}:\ |\mathcal{T}f(z)|>\lambda\}\big|\leq C\,\lambda^{-q}\|f\|_{L^{p}({\mathbb{D}})}^{\,q},$
with the standard modification when $q=\infty$.
Moreover, we say that $\mathcal{T}$ is of restricted weak type $(p,q)$ if $\mathcal{T}:L^{p,1}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded, or equivalently (a well-known fact), if there exists $C>0$ such that for every measurable set $E\subset{\mathbb{D}}$ with $|E|<\infty$ and all $\lambda>0$,
$\big|\{z\in{\mathbb{D}}:\ |\mathcal{T}\mbox{1\hskip-4.25pt{1}}_{E}(z)|>\lambda\}\big|\leq C\,\lambda^{-q}|E|^{q/p},$
again with the usual modification when $q=\infty$.
Finally, given a nonnegative locally integrable function $\omega$ on $\mathbb{D}$}, referred to as a weight, we define the weighted space $L^{p}({\mathbb{D}},\omega)$ to be the collection of all measurable functions on $\mathbb{D}$} satisfying $\|f\|^{p}_{L^{p}({\mathbb{D}},\omega)}:=\int_{{\mathbb{D}}}|f(z)|^{p}\,\omega(z)\,dz<+\infty$, with the usual modification when $p=\infty$. A weight $\omega$ is called radial if $\omega(z)$ depends only on $|z|$, that is, $\omega(z)=\omega(|z|)$ for all $z\in{\mathbb{D}}$.
3. $L^{p}$ theory for the hypersingular maximal operator
In this section, we divide our analysis of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ into two distinct regimes:
  1. (1)
    Off-critical line regime, that is, when $(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\}$;
  2. (2)
    Critical line regime, that is, when $(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\}$.
Here and henceforth, we always assume that $1<t<3/2$.
3.1. Off-critical line estimate.
We begin with the following observation.
Observation 3.1.
For any $\varepsilon>0$, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\infty}({\mathbb{D}})\to L^{\frac{1}{2t-2+\varepsilon}}({\mathbb{D}})$ is bounded.
Proof.
The proof follows from a direct computation. Indeed,
$\int_{{\mathbb{D}}}\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{2t-2+\varepsilon}}dA(z)$ $\lesssim\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}\int_{{\mathbb{D}}}\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}1(z)\right|^{\frac{1}{2t-2+\varepsilon}}dA(z)$
$\simeq\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2(t-1)+\varepsilon}}}dA(z)$
$\simeq\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}\int_{0}^{1}\frac{dr}{(1-r)^{\frac{2(t-1)}{2(t-1)+\varepsilon}}}\lesssim\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}.$
From the above arguments, it is not hard to see that ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ is unbounded from $L^{\infty}({\mathbb{D}})$ to $L^{\frac{1}{2t-2}}({\mathbb{D}})$. ∎
Next, we prove estimate near the endpoint $\left(1/p,1/q\right)=(3-2t,1)$.
Lemma 3.2.
For any $0<\varepsilon\leq 3-2t$, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})\to L^{1}({\mathbb{D}})$ is bounded.
Proof.
Without loss of generality, we may assume $\varepsilon<3-2t$. The case when $\varepsilon=3-2t$ is obvious from the argument in (1.2).
Let $\alpha>0$ and denote $E_{\alpha}:=\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\right\}$. Observe that one can decompose $E_{\alpha}$ into a collection of maximal and mutually disjoint Carleson boxes $\{Q_{\alpha,i}\}_{i\geq 1}$ such that
(3.1) $\frac{1}{|Q_{\alpha,i}|^{t}}\int_{Q_{\alpha,i}}|f(z)|dA(z)>\alpha,\qquad\textnormal{for each}\;\;i\geq 1.$
Therefore,
$\alpha|E_{\alpha}|$ $=\alpha\sum_{i=1}^{\infty}|Q_{\alpha,i}|\leq\sum_{i=1}^{\infty}\frac{|Q_{\alpha,i}|}{|Q_{\alpha,i}|^{t-1}}\cdot\frac{1}{|Q_{\alpha,i}|}\int_{Q_{\alpha,i}}|f(z)|dA(z)$
$\leq\sum_{i=1}^{\infty}\frac{1}{|Q_{\alpha,i}|^{t-1}}\int_{Q_{\alpha,i}^{\textnormal{up}}}{\mathcal{M}}^{{\mathcal{D}}}f(z)dA(z)$
$\leq\sum_{i=1}^{\infty}\frac{1}{|Q_{\alpha,i}|^{t-1}}\left(\int_{Q^{\textnormal{up}}_{\alpha,i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\left(\int_{Q_{\alpha,i}}dA(z)\right)^{2t+\varepsilon-2}$
$=\sum_{i=1}^{\infty}|Q_{\alpha,i}|^{t-1+\varepsilon}\left(\int_{Q^{\textnormal{up}}_{\alpha,i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}$
(3.2) $\leq\left(\sum_{i=1}^{\infty}\int_{Q^{\textnormal{up}}_{\alpha,i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{i=1}^{\infty}|Q_{\alpha,i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2},$
where ${\mathcal{M}}^{{\mathcal{D}}}={\mathcal{M}}_{1}^{{\mathcal{D}}}$ is the standard maximal operator over dyadic Carleson tents, and in the last estimate, we have used the fact that $\{Q_{\alpha,i}\}_{i\geq 1}$ are mutually disjoint.
Now for any $f\in L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})$, we have to estimate $\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f\right\|_{L^{1}({\mathbb{D}})}$. Using (3.1), we have
$\int_{{\mathbb{D}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)$ $dA(z)=\sum_{k\in\mathbb{Z}}\int_{E_{4^{tk}}\backslash E_{4^{t(k+1)}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)dA(z)\lesssim\sum_{k\in\mathbb{Z}}4^{tk}\left|E_{4^{tk}}\right|$
$\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{Q^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{i=1}^{\infty}|Q_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
(3.3) $\leq\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|Q_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
Since for each $k\in\mathbb{Z}$, $\{Q_{4^{tk},i}\}_{i\geq 1}$ is maximal, we have $\left\{Q^{\textnormal{up}}_{4^{tk},i}\right\}_{k\in\mathbb{Z},i\geq 0}$ are mutually disjoint, therefore, the first double sum in (3.1) is bounded above by
(3.4) $\int_{{\mathbb{D}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)=\left\|{\mathcal{M}}^{{\mathcal{D}}}f\right\|^{\frac{1}{3-2t-\varepsilon}}_{L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})}\lesssim\left\|f\right\|^{\frac{1}{3-2t-\varepsilon}}_{L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})},$
where in the last estimate above, we have used the boundedness of ${\mathcal{M}}^{{\mathcal{D}}}:L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})\to L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})$.
We are left with estimating the second double summation in (3.1). Indeed, we have
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|Q_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}$ $\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|Q^{\textnormal{up}}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}=\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left|Q^{\textnormal{up}}_{4^{tk},i}\right|\cdot\frac{1}{\left|Q^{\textnormal{up}}_{4^{tk},i}\right|^{\frac{t-1}{2t+\varepsilon-2}}}$
$\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q^{\textnormal{up}}_{4^{tk},i}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)$
$=\int_{0}^{1}\frac{1}{(1-r)^{\frac{2(t-1)}{2t+\varepsilon-2}}}dr<+\infty.$
The desired claim follows by combining the above estimate with (3.1) and (3.4). ∎
Therefore, we derive the following result.
Proposition 3.3.
For any $(p,q)$ belonging to the off-critical line regime, namely,
$(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\},$
one has ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q}({\mathbb{D}})$ is bounded.
Proof.
The desired claim follows clearly by interpolating the estimates derived in Observation 3.1 and Lemma 3.2. ∎
3.2. Critical line estimate
We first deal with the endpoint $\left(\frac{1}{p},\frac{1}{q}\right)=(0,2t-2)$.
Lemma 3.4.
${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\infty}({\mathbb{D}})\to L^{\frac{1}{2t-2},\infty}({\mathbb{D}})$ is bounded.
Proof.
Take any $\alpha>0$ and any measurable function $f$ with $\left\|f\right\|_{L^{\infty}({\mathbb{D}})}=1$. Since $\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\right|>\alpha$, we have
$\alpha\leq\sup_{z\in Q_{I},\;I\in{\mathcal{D}}}\frac{1}{|Q_{I}|^{t-1}}\simeq\frac{1}{(1-|z|^{2})^{2(t-1)}}.$
Therefore,
$\left|\left\{z\in{\mathbb{D}}:\;\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\right|>\alpha\right\}\right|$ $\leq\left|\left\{z\in{\mathbb{D}}:\sup_{z\in Q_{I},\;I\in{\mathcal{D}}}\frac{1}{|Q_{I}|}>\alpha^{\frac{1}{t-1}}\right\}\right|$
$\lesssim\left|\left\{z\in{\mathbb{D}}:\;\frac{1}{(1-|z|^{2})^{2}}>\alpha^{\frac{1}{t-1}}\right\}\right|$
$=\left|\left\{z\in{\mathbb{D}}:1-\alpha^{-\frac{1}{2(t-1)}}\leq|z|^{2}<1\right\}\right|$
$\simeq\alpha^{-\frac{1}{2(t-1)}},$
which gives
$\alpha\left|\left\{z\in{\mathbb{D}}:\;\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\right|>\alpha\right\}\right|^{2t-2}\lesssim 1.$
The proof is complete. ∎
Next, we treat the other endpoint $\left(\frac{1}{p},\frac{1}{q}\right)=(3-2t,1)$.
Theorem 3.5.
Let $\omega$ be a weight on $[0,1)$ satisfying $\omega(r)\geq c>0$ for all $r\in[0,1/2)$, and let $\omega(z):=\omega(|z|)$ denote the associated radial weight on $\mathbb{D}$}. Then ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)\to L^{1,\infty}({\mathbb{D}})$ is bounded if and only if
(3.5) $\sup_{k\geq 0}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{1}{\omega^{\frac{3-2t}{2t-2}}(r)}dr<+\infty.$
Proof.
Sufficiency. Let $\alpha>0$ and denote $E:=\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\right\}$. As usual, we decompose $E$ into a union of maximal and mutually disjoint Carleson boxes $\{Q_{i}\}_{i\geq 1}$ satisfying $Q_{i}=Q_{I_{i}}$ for $I_{i}\subset\mathbb{T}$ and
$\frac{1}{|Q_{i}|^{t}}\int_{Q_{i}}|f(z)|dA(z)>\alpha\qquad\textnormal{for each}\;\;i\geq 1.$
This gives
$\alpha|E|$ $=\alpha\sum_{i=1}^{\infty}|Q_{i}|\leq\sum_{i=1}^{\infty}\frac{1}{|Q_{i}|^{t-1}}\int_{Q_{i}}|f(z)|dA(z)$
$=\sum_{i=1}^{\infty}\frac{1}{|Q_{i}|^{t-1}}\left(\sum_{J\subseteq I_{i},\;J\;\textnormal{dyadic}}\int_{Q_{J}^{\textnormal{up}}}|f(z)|\omega^{3-2t}(z)\cdot\frac{1}{\omega^{3-2t}(z)}dA(z)\right)$
$\leq\sum_{i=1}^{\infty}\frac{1}{|Q_{i}|^{t-1}}\sum_{J\subseteq I_{i},\;J\;\textnormal{dyadic}}\left(\int_{Q_{J}^{\textnormal{up}}}|f(z)|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
$\leq\left(\sum_{i=1}^{\infty}\sum_{J\subseteq I_{i},\;J\;\textnormal{dyadic}}\int_{Q_{J}^{\textnormal{up}}}|f(z)|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\sum_{J\subseteq I_{i},\;J\;\textrm{dyadic}}\frac{1}{|Q_{i}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
(3.6) $\leq\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}\cdot\left(\sum_{i=1}^{\infty}\sum_{J\subseteq I_{i},\;J\;\textrm{dyadic}}\frac{1}{|Q_{i}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
For the double summation in the above estimate, we have [^3] Here and henceforth, we normalize the length of $\mathbb{T}$ so that $|\mathbb{T}|=1$. In particular, any dyadic descendant of $\mathbb{T}$ has length $2^{-k}$ for some $k\geq 0$.
$\sum_{i=1}^{\infty}\sum_{J\subseteq I_{i},\;J\;\textrm{dyadic}}\frac{1}{|Q_{i}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}$ $\simeq\sum_{i=1}^{\infty}\sum_{k\geq 0}\sum_{\begin{subarray}{c}J\subseteq I_{i},\;J\;\textrm{dyadic}\\ |J|=2^{-k}|I_{i}|\end{subarray}}\frac{1}{2^{k}|Q_{J}^{\textnormal{up}}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}$
(3.7) $\simeq\sum_{k\geq 0}\frac{1}{2^{k}}\left(\sum_{i=1}^{\infty}\sum_{\begin{subarray}{c}J\subseteq I_{i},\;J\;\textrm{dyadic}\\ |J|=2^{-k}|I_{i}|\end{subarray}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right).$
Using now the assumption that $\omega$ is radial and (3.5), we have
$\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}$ $\simeq\frac{1}{|J|}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}\lesssim|J|\cdot\frac{1}{|J|}\int_{1-|J|}^{1-\frac{|J|}{2}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\lesssim|J|.$
Substituting the above estimate back to (3.2), we have
RHS of (3.2) $\lesssim\sum_{k\geq 0}\frac{1}{2^{k}}\left(\sum_{i=1}^{\infty}\sum_{\begin{subarray}{c}J\subseteq I_{i},\;J\;\textrm{dyadic}\\ |J|=2^{-k}|I_{i}|\end{subarray}}|J|\right)$
$=\sum_{k\geq 0}\frac{1}{2^{k}}\left(\sum_{i=1}^{\infty}|I_{i}|\right)<+\infty,$
where in the last estimate we have used the fact that $I_{i}$’s are mutually disjoint. The desired weak-type estimate ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)\to L^{1,\infty}({\mathbb{D}})$ then follows by plugging the above estimate back into (3.2).
Necessity. For each $k\geq 0$, denote
(3.8) $D_{k}:=\left\{z\in{\mathbb{D}}:1-\frac{1}{2^{k}}\leq|z|<1-\frac{1}{2^{k+1}}\right\},$
and $f_{k}(z):=\omega^{-\frac{3-2t}{2t-2}}(z)\mbox{1\hskip-4.25pt{1}}_{D_{k}}(z)$. A direct computation yields
(3.9) $\left\|f_{k}\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}=\left(\int_{D_{k}}\omega^{-\frac{3-2t}{2t-2}}(z)dA(z)\right)^{3-2t}\simeq\left(\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right)^{3-2t}.$
Next, consider ${\mathcal{D}}_{k}:=\left\{I\in{\mathcal{D}}:|I|=2^{-k}\right\}$. Then for each $w\in Q_{I}$ with $I\in{\mathcal{D}}_{k}$, one has
${\mathcal{M}}_{t}^{{\mathcal{D}}}f_{k}(w)$ $\geq\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}}f_{k}(z)dA(z)\geq\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}^{\textnormal{up}}}\frac{1}{\omega^{\frac{3-2t}{2t-2}}(z)}dA(z)$
$\simeq\frac{1}{|Q_{I}|^{t}}\cdot|I|\cdot\int^{1-\frac{|I|}{2}}_{1-|I|}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\simeq 2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)},$
which implies
$Q_{I}\subseteq\left\{w\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{k}(w)>\widetilde{c}2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right\}$
for some absolute constant $\widetilde{c}>0$ being sufficiently small. Since $I\in{\mathcal{D}}_{k}$ are mutually disjoint, we further get
(3.10) $\bigcup_{I\in{\mathcal{D}}_{k}}Q_{I}\subseteq\left\{w\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{k}(w)>\widetilde{c}2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right\},$
Therefore, by the assumption ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)\to L^{1,\infty}({\mathbb{D}})$ and (3.9), we have
$2^{-k}$ $\simeq\left|\bigcup_{I\in{\mathcal{D}}_{k}}Q_{I}\right|\lesssim\left|\left\{w\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{k}(w)>c2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right\}\right|$
$\lesssim\frac{1}{2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}}\cdot\left\|f_{k}\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}$
$\lesssim\frac{1}{2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}}\left(\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right)^{3-2t}$
$=\frac{1}{2^{k(2t-1)}}\cdot\left(\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right)^{2-2t},$
which gives
$2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\lesssim 1.$
The proof of the necessity is complete. ∎
Remark 3.6.
In Theorem 3.5, the assumption $\omega(r)\geq c>0$ for $r\in\left[0,\frac{1}{2}\right)$ is used only to ensure that $\int_{0}^{1/2}\omega(r)^{-\frac{3-2t}{2t-2}}\,dr<\infty$. This is a minor technical requirement, since the relevant (and more delicate) behavior of the weight occurs near the boundary, as $r\to 1^{-}$.
As a direct application of Theorem 3.5 with $\omega\equiv 1$, we have ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t}}({\mathbb{D}})\to L^{1,\infty}$ is bounded. Interpolating this with Lemma 3.4, we derive the following estimates on the critical line.
Corollary 3.7.
For any $(p,q)$ belonging to the critical line regime, namely,
$(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in(0,1)^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\},$
then for every $0<r\leq\infty$, the operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ extends a bounded operator from $L^{p,r}({\mathbb{D}})$ to $L^{q,r}({\mathbb{D}})$. In particular, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
Proof.
The first assertion follows from the off-diagonal Marcinkiewicz interpolation theorem [14, Theorem 1.4.19]. The second follows from by letting $r=\infty$ and the fact that $L^{p}({\mathbb{D}})\subset L^{p,\infty}({\mathbb{D}})$. ∎
A natural question arising from Theorem 3.5 is whether ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t}}({\mathbb{D}})\to L^{1}({\mathbb{D}})$ is bounded or not. Here, again, we would like to formulate this result in the setting of radial weights. For this purpose, we recall the notion of Békollé–Bonami weights. Let $1<l<\infty$. We say that a weight $\omega$ on $\mathbb{D}$} belongs to the Békollé–Bonami class ${\bf B}_{l}$ if
(3.11) $\omega]_{{\bf B}_{l}}:=\sup_{I\subseteq\mathbb{T}}\left(\frac{1}{|Q_{I}|}\int_{Q_{I}}\omega\,dA\right)\left(\frac{1}{|Q_{I}|}\int_{Q_{I}}\omega^{-\frac{1}{l-1}}\,dA\right)^{l-1}<\infty,$
In particular, if $\omega$ is radial, then the above condition is equivalent to
$\omega]_{{\bf B}_{l}}\simeq\sup_{0<h<1}\left(\frac{1}{h}\int_{1-h}^{1}\omega(r)\,dr\right)\left(\frac{1}{h}\int_{1-h}^{1}\omega(r)^{-\frac{1}{l-1}}\,dr\right)^{l-1}<+\infty$
An important reason to consider the Békollé–Bonami weights is that they provide a necessary and sufficient condition for $\mathcal{M}$} to extend to a bounded operator on $L^{l}({\mathbb{D}},\omega)$, with operator norm of magnitude $\omega]_{{\bf B}_{l}}^{\frac{1}{l-1}}$ (see, e.g., [1]). Here, $\mathcal{M}$} refers to the standard maximal operator over all Carleson tents.
Theorem 3.8.
Let $\omega$ be a radial weight that satisfies the assumption of Theorem 3.5. Let further, $\omega\in{\bf B}_{\frac{1}{3-2t}}$. Then ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)\to L^{1}({\mathbb{D}})$ is bounded if and only if
$\sum_{k\geq 0}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{1}{\omega^{\frac{3-2t}{2t-2}}(r)}dr<+\infty.$
In particular, ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ maps $L^{\frac{1}{3-2t}}({\mathbb{D}})$ unboundedly into $L^{1}({\mathbb{D}})$.
Proof.
We prove Theorem 3.8 by adapting the ideas from the proof of Lemma 3.2 and Theorem 3.5.
Sufficiency. Again, for any $\alpha>0$, we denote $E_{\alpha}:=\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\right\}$. Then as usual, one can decompose $E_{\alpha}$ into disjoint union of maximal Carleson boxes $\{Q_{\alpha,i}\}_{i\geq 1}$ enjoying (3.1). Then following the argument in (3.1) with applying Hölder for the conjugate pair $(1/l,1/l^{\prime})=(3-2t,2t-2)$, we have
$\alpha|E_{\alpha}|$ $\lesssim\sum_{i=1}^{\infty}\frac{1}{|Q_{\alpha,i}|^{t-1}}\int_{Q_{\alpha,i}^{\textnormal{up}}}{\mathcal{M}}^{{\mathcal{D}}}f(z)\omega^{3-2t}(z)\cdot\frac{1}{\omega^{3-2t}(z)}dA(z)$
$\lesssim\left(\sum_{i=1}^{\infty}\int_{Q_{\alpha,i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{Q_{\alpha,i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}.$
Therefore,
$\int_{{\mathbb{D}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)$ $dA(z)=\sum_{k\in\mathbb{Z}}\int_{E_{4^{tk}}\backslash E_{4^{t(k+1)}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)dA(z)\lesssim\sum_{k\in\mathbb{Z}}4^{tk}\left|E_{4^{tk}}\right|$
$\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
$\lesssim\left\|{\mathcal{M}}^{{\mathcal{D}}}f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
(3.12) $\lesssim[\omega]_{{\bf B}_{\frac{1}{3-2t}}}^{\frac{3-2t}{2t-2}}\cdot\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2},$
where in the last estimate, we used the assumption $\omega\in{\bf B}_{\frac{1}{3-2t}}$. Finally for the double summation in (3.2), using the fact that $\{Q_{4^{tk},i}^{\textnormal{up}}\}_{k\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, we have
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}$ $\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}dA(z)$
$\simeq\int_{0}^{1}\frac{rdr}{(1-r^{2})\omega^{\frac{3-2t}{2t-2}}(r)}$
$\simeq\sum_{k\geq 0}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{1}{\omega^{\frac{3-2t}{2t-2}}(r)}dr<+\infty.$
The sufficiency is therefore proved.
Necessity. Assume ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)\to L^{1}({\mathbb{D}})$ is bounded. Let $N\in{\mathbb{N}}$ be sufficiently large, and take the test function
$f_{N}(z):=\sum_{k=0}^{N}2^{k(3-2t)}\omega^{-\frac{3t-2}{2t-2}}(z)\mbox{1\hskip-4.25pt{1}}_{D_{k}}(z),$
where $D_{k}$ is the annulus defined as in (3.8).
On one hand side, recall that $\{D_{k}\}_{k\geq 0}$ are mutually disjoint, we have
$\left\|f_{N}\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}$ $=\left(\int_{\mathbb{D}}\left|\sum_{k=0}^{N}2^{k(3-2t)}\omega^{-\frac{3t-2}{2t-2}}(z)\mbox{1\hskip-4.25pt{1}}_{D_{k}}(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}$
$=\left(\sum_{k^{\prime}\geq 0}\int_{D_{k}^{\prime}}\left|\sum_{k=0}^{N}2^{k(3-2t)}\omega^{-\frac{3t-2}{2t-2}}(z)\mbox{1\hskip-4.25pt{1}}_{D_{k}}(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}$
$=\left(\sum_{k=0}^{N}\int_{D_{k}}\left|2^{k(3-2t)}\omega^{-\frac{3t-2}{2t-2}}(z)\mbox{1\hskip-4.25pt{1}}_{D_{k}}(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}$
(3.13) $=\left(\sum_{k=0}^{N}2^{k}\int_{D_{k}}\omega^{-\frac{3-2t}{2t-2}}(z)dA(z)\right)^{3-2t}\simeq\left(\sum_{k=0}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right)^{3-2t}.$
On the other hand side, by the argument in (3.10), we see that for $1\leq k\leq N$,
${\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}(z)\gtrsim 2^{k(3-2t)}\cdot 2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}=2^{2k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)},$
and hence
$\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}\right\|_{L^{1}({\mathbb{D}})}$ $\geq\sum_{k=1}^{N}\int_{D_{k}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}(z)dA(z)$
$\gtrsim\sum_{k=1}^{N}2^{-k}\cdot 2^{2k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}$
(3.14) $=\sum_{k=1}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}.$
Finally, since $\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}\right\|_{L^{1}({\mathbb{D}})}\lesssim\left\|f_{N}\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}$, using this with (3.2) and (3.2), we derive that
(3.15) $\sum_{k=0}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\lesssim\left(\sum_{k=0}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right)^{3-2t}.$
Note that since $\omega$ is a weight (hence locally integrable), this means
$\sum_{k=0}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\simeq\int_{0}^{1-\frac{1}{2^{N+1}}}\frac{dr}{(1-r)\omega^{\frac{3-2t}{2t-2}}(r)}<+\infty,$
and hence (3.15) gives
$\left(\sum_{k=0}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right)^{2t-2}\lesssim 1,$
where the implicit constant in the above estimate is independent of the choice of $N$. Finally, the desired necessary part follows by letting $N\to\infty$ in the above estimate and the assumption that $1<t<3/2$. ∎
4. Critical line estimates for the hypersingular Bergman projection
Our next goal is to establish critical line estimates for the hypersingular Bergman projection: for $1<t<3/2$,
$K_{2t}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}dA(w),$
which can be regarded as a singular integral counterpart of the hypersingular maximal operator $M_{t}^{{\mathcal{D}}}$. We first recall that the off-critical line $L^{p}$ theory was studied in [12, Theorem 3]. In particular, they showed that for $K_{2t}$ is bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ if and only 1/q1/p>2t21/q-1/p>2t-2 for $1\leq p,q\leq\infty$, which is exactly the off-critical line regime for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$. Therefore, it is natural to consider the behavior of $K_{2t}$ on the critical line
$\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\}.$
We first have the following observation.
Lemma 4.1.
For any $1<t<3/2$, $K_{2t}:L^{\infty}({\mathbb{D}})\to L^{\frac{1}{2t-2},\infty}({\mathbb{D}})$ is bounded.
Proof.
The proof of this lemma is straightforward. Let $f\in L^{\infty}({\mathbb{D}})$ with $\left\|f\right\|_{L^{\infty}({\mathbb{D}})}=1$. Observe now that
$\left|K_{2t}f(z)\right|=\left|\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}dA(w)\right|\lesssim\int_{{\mathbb{D}}}\frac{1}{|1-z\overline{w}|^{2t}}dA(w)\simeq\frac{1}{(1-|z|^{2})^{2(t-1)}},$
where in the last estimate, we used the standard integral estimate [31, Theorem 1.12]. Therefore, for any $\alpha>0$, as in Lemma 3.4,
$\left|\left\{z\in{\mathbb{D}}:\left|K_{2t}f(z)\right|>\alpha\right\}\right|\lesssim\alpha^{-\frac{1}{2(t-1)}},$
which yields the desired weak-type bounds. ∎
Our next goal is to establish estimates at the other endpoint $\big(\tfrac{1}{3-2t},1\big)$. First, we observe the following unboundedness result.
Proposition 4.2.
Let $1<t<3/2$. Then the hypersingular Berezin transform $K_{2t}^{+}$ fail to be bounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$, where we recall that
$K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w).$
Proof.
Take any $\varepsilon>0$ sufficiently small such that $3-2t-\varepsilon>0$, and let $f(z):=\frac{1}{(1-|z|^{2})^{3-2t-\varepsilon}},\;z\in{\mathbb{D}}$. First, by a direct computation, we have
$\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}})}\simeq\frac{1}{\varepsilon^{3-2t}}.$
On the other hand, by [31, Theorem 1.12], one has
$K_{2t}^{+}f(z)$ $=\int_{{\mathbb{D}}}\frac{1}{(1-|w|^{2})^{3-2t-\varepsilon}|1-z\overline{w}|^{2t}}dA(w)$
$=\int_{{\mathbb{D}}}\frac{(1-|w|^{2})^{2t+\varepsilon-3}}{|1-z\overline{w}|^{2t}}dA(w)$
$\simeq\frac{1}{(1-|z|^{2})^{1-\varepsilon}}.$
Therefore, for any $\alpha>0$, the above estimate yields
$\alpha\left|\left\{z\in{\mathbb{D}}:\left|K_{2t}^{+}f(z)\right|\gtrsim\alpha\right\}\right|$ $\simeq\alpha\left|\left\{z\in{\mathbb{D}}:\frac{1}{(1-|z|^{2})^{1-\varepsilon}}>\alpha\right\}\right|$
$\simeq\alpha\cdot\alpha^{-\frac{1}{1-\varepsilon}}=\alpha^{\frac{\varepsilon}{1-\varepsilon}}.$
Thus, if $K_{2t}^{+}:L^{\frac{1}{3-2t}}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded, then for any $\alpha>0$, one has
$\alpha^{\frac{\varepsilon}{1-\varepsilon}}\lesssim\frac{1}{\varepsilon^{3-2t}},$
which is impossible by letting $\alpha\to\infty$. ∎
Now we turn to some positive results. As discussed in Section 1 (see the comments following (1.3)), the main difficulty is the lack of strong-type bounds along the critical line 1/q1/p=2t21/q-1/p=2t-2. To overcome this difficulty, we combine a “hypersingular” sparse domination with Bourgain’s interpolation trick. (see, e.g., [3, 6]; see also [7] for an abstract extension of this method developed in the setting of fairly general normed vector spaces). The following formulation is the one needed for our purposes.
Lemma 4.3 (Bourgain’s interpolation trick).
Let $\beta_{1},\beta_{2}>0$ and $\{T_{j}\}_{j\geq 0}$ be a collection of sublinear operators satisfying
$\left\|T_{j}\right\|_{L^{p_{1}}({\mathbb{D}})\to L^{q_{1}}({\mathbb{D}})}\leq M_{1}2^{\beta_{1}j}$
and
$\left\|T_{j}\right\|_{L^{p_{2}}({\mathbb{D}})\to L^{q_{2}}({\mathbb{D}})}\leq M_{2}2^{-\beta_{2}j},$
for some $1\leq p_{1},p_{2},q_{1},q_{2}\leq+\infty$ and $M_{1},M_{2}>0$, then $T=\sum_{j\geq 1}T_{j}$ enjoys restricted weak type estimate between the intermediate spaces:
$\left\|T\right\|_{L^{p,1}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})}\leq CM_{1}^{\theta}M_{2}^{1-\theta},$
where
$\theta=\frac{\beta_{2}}{\beta_{1}+\beta_{2}},\quad\frac{1}{p}=\frac{\theta}{p_{1}}+\frac{1-\theta}{p_{2}},\quad\textrm{and}\quad\frac{1}{q}=\frac{\theta}{q_{1}}+\frac{1-\theta}{q_{2}},$
and $C$ depends only on $\beta_{1}$ and $\beta_{2}$.
We are ready to state the main result in this section.
Theorem 4.4.
For any $1<t<3/2$, $K_{2t}:L^{\frac{1}{3-2t},1}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded.
Proof.
Let $f$ be any measurable function on $\mathbb{D}$}. We first observe that $K_{2t}f$ admits a pointwise bound by a finite sum of “hypersingular” sparse operators. This follows from standard dyadic analysis on $\mathbb{D}$} (see, e.g., [13, 16, 25]). For the reader’s convenience, we include a brief sketch of the argument.
First, recall that for any $z,w\in{\mathbb{D}}$, there exists a Carleson tent $Q_{J}$, containing both $z$ and $w$ such that $|Q_{J}|\simeq\left|1-z\overline{w}\right|^{2}$ (see, e.g., [25, Section 3]). Then using the $1/3$-trick, one can find some $I\in{\mathcal{D}}$ or $\widetilde{{\mathcal{D}}}$ (see, Remark 1.4), such that
$J\subset I\quad\textrm{and}\quad|I|\lesssim|J|,$
which further gives
$z,w\in Q_{I}\quad\textrm{and}\quad|Q_{I}|\simeq|1-z\overline{w}|^{2}.$
As a conseqeunce, we have the following pointwise “hypersingular” sparse bound
(4.1) $\left|K_{2t}f(z)\right|=\left|\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}dA(w)\right|\lesssim{\mathcal{A}}^{t}_{{\mathcal{S}},{\mathcal{D}}}f(z)+{\mathcal{A}}^{t}_{{\mathcal{S}},\widetilde{{\mathcal{D}}}}f(z),\quad z\in{\mathbb{D}},$
where ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ is the hypersingular sparse operator associated to a dyadic system $\mathcal{D}$} on $\mathbb{T}$, defined as [^4] We call ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ a hypersingular sparse operator since, on the critical line $\{(\tfrac{1}{p},\tfrac{1}{q}):\ \tfrac{1}{q}-\tfrac{1}{p}=2t-2\}$, it fails to satisfy strong $(p,q)$ bounds. Indeed, any strong-type estimate for ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ at a point on this line would imply the corresponding bound for $K_{2t}$, which is known to be false by [12, Theorem 3].
${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}f(z):=\sum_{I\in{\mathcal{D}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q_{I}}(z)}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|dA(w).$
Therefore, it suffices to prove that $A_{{\mathcal{S}},{\mathcal{D}}}^{t}:L^{\frac{1}{3-2t},1}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded. For simplicity, we denote ${\mathcal{A}}_{{\mathcal{S}}}^{t}={\mathcal{A}}^{t}_{{\mathcal{S}},{\mathcal{D}}}$.
Now we shall apply Bourgain’s interpolation trick. To do this, we decompose
${\mathcal{A}}^{t}_{{\mathcal{S}}}=\sum_{j\geq 0}{\mathcal{A}}^{t}_{{\mathcal{S}},j},$
where
${\mathcal{A}}^{t}_{{\mathcal{S}},j}f(z):=\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q_{I}}(z)}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|dA(w),\quad z\in{\mathbb{D}}.$
We claim that for any $j\geq 0$,
(4.2) $\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}\right\|_{L^{1}({\mathbb{D}})\to L^{1}({\mathbb{D}})}\lesssim 2^{(2t-2)j}$
and
(4.3) $\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}\right\|_{L^{\infty}({\mathbb{D}})\to L^{1}({\mathbb{D}})}\lesssim 2^{-(3-2t)j}.$
For the moment, we assume that (4.2) and (4.3) hold. To see the desired restrict weak type bound for ${\mathcal{A}}_{{\mathcal{S}}}^{t}$, we apply Lemma 4.3 with
$\beta_{1}=2t-2,\quad\beta_{2}=3-2t,\quad p_{1}=q_{1}=q_{2}=1,\quad\textrm{and}\quad p_{2}=\infty.$
Therefore,
$\left\|{\mathcal{A}}_{{\mathcal{S}}}^{t}\right\|_{L^{p,1}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})}\lesssim 1,$
for
$\theta=3-2t,\quad\frac{1}{p}=\frac{3-2t}{1}+\frac{2t-2}{p_{2}}=3-2t,\quad\textrm{and}\quad\frac{1}{q}=1.$
This concludes the proof.
Therefore, we only need to show (4.2) and (4.3).
Proof of (4.2). Note that it is more convenient to work with ${\mathcal{A}}_{{\mathcal{S}},j}^{t}$ rather than ${\mathcal{A}}_{{\mathcal{S}}}^{t}$, since the family $\{Q_{I}\}_{I\in{\mathcal{D}},\ |I|=2^{-j}}$ is pairwise disjoint. The argument is straightforward: let $f\in L^{1}({\mathbb{D}})$,
$\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}f\right\|_{L^{1}({\mathbb{D}})}$ $=\int_{{\mathbb{D}}}\left(\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q_{I}}(z)}{|Q_{I}|^{t}}\int_{Q_{I}}|f(w)|dA(w)\right)dA(z)$
$=\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\frac{1}{|Q_{I}|^{t-1}}\int_{Q_{I}}|f(w)|dA(w)$
$\simeq 2^{(2t-2)j}\sum_{\begin{subarray}{c}I\in{\mathcal{D}}\\ |I|=2^{-j}\end{subarray}}\int_{Q_{I}}|f(w)|dA(w)$
$=2^{(2t-2)j}\int_{1-2^{-j}\leq|z|<1}|f(w)|dA(w)$
$\leq 2^{(2t-2)j}\left\|f\right\|_{L^{1}({\mathbb{D}})},$
which gives (4.2).
Proof of (4.3). The proof for the second claim is similar. Indeed, for any $f\in L^{\infty}({\mathbb{D}})$, from the above argument, we have
$\left\|{\mathcal{A}}_{{\mathcal{S}},j}^{t}f\right\|_{L^{1}({\mathbb{D}})}$ $\simeq 2^{(2t-2)j}\int_{1-2^{-j}\leq|z|<1}|f(w)|dA(w)$
$\lesssim 2^{(2t-2)j}\cdot\left|\left\{z\in{\mathbb{D}}:1-2^{-j}\leq|z|<1\right\}\right|\cdot\left\|f\right\|_{L^{\infty}({\mathbb{D}})}$
$\simeq 2^{(2t-2)j}\cdot 2^{-j}\cdot\left\|f\right\|_{L^{\infty}({\mathbb{D}})}$
$=2^{-(3-2t)j}\left\|f\right\|_{L^{\infty}({\mathbb{D}})}.$
Hence, (4.3) holds.
The proof of Theorem 4.4 is complete. ∎
Combining Lemma 4.1 and Theorem 4.4 with off-diagonal Marcinkiewicz interpolation theorem, we have the following.
Corollary 4.5.
For any $(p,q)$ belonging to the critical line regime, that is,
$\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in(0,1)^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\},$
then for every $0<r\leq\infty$, the operator $K_{2t}^{{\mathcal{D}}}$ extends to a bounded map
$K_{2t}^{{\mathcal{D}}}:L^{p,r}({\mathbb{D}})\to L^{q,r}({\mathbb{D}}),$
In particular, $K_{2t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
Remark 4.6.
Using Bourgain’s interpolation trick, one can give an alternative proof of Corollary 3.7. Indeed, the following pointwise bound is immediate:
${\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\leq{\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}f(z),\quad z\in{\mathbb{D}}.$
The desired conclusion then follows by repeating the proof of Theorem 4.4 and applying the off-diagonal Marcinkiewicz interpolation theorem.
5. Hypersingular sparse operators
A natural question arising from the proof of Theorem 4.4 is whether Bourgain’s interpolation trick can be used to study other dyadic operators in harmonic analysis. It turns out that a more general principle underlies this phenomenon.
Let us now turn to some details. We recall some definitions first.
Definition 5.1.
Let $\mathcal{D}$} be a dyadic system in $\mathbb{R}^{n}$ and ${\mathcal{S}}\subset{\mathcal{D}}$ be a collection of dyadic cubes. For $0<\eta<1$, we say $\mathcal{S}$} is $\eta$-sparse, if there exists a collection of measurable sets $\{E(Q)\}_{Q\in{\mathcal{S}}}$, such that
  1. (1)
    $E(Q)\subseteq Q$ for all $Q\in{\mathcal{S}}$;
  2. (2)
    $|E(Q)|\geq\eta\mu(Q)$ for all $Q\in{\mathcal{S}}$;
  3. (3)
    $E(Q)\cap Q^{\prime}=\emptyset$ for any $^{\prime}\subset Q, QQ$$^{\prime}\in Q$.
Remark 5.2.
We remark that the above notion of sparse families was introduced by Lerner, Lorist, and Ombrosi [18] in their work on operator-free sparse domination. Compared with the usual definition used in the literature, this formulation is slightly more restrictive, most through the third condition above, which we refer to as the contracting property. A simple observation shows that this contracting property forces the sets $\{E(Q)\}_{Q\in\mathcal{S}}$ to be pairwise disjoint, thereby recovering the usual disjointness requirement in the standard definition of a sparse family.
Such sparse collections arise naturally in many applications of sparse domination, for instance: (1) collections of Carleson boxes; and (2) sparse collections produced by stopping-time constructions, such as those associated with maximal operators and Calderón–Zygmund operators. We refer the reader to [18] for further discussion and applications of sparse families with the contracting property.
We begin with the following model case. Let $t>1$, and for simplicity, assume that $\mathcal{S}$ be an $\eta$-sparse family of cubes in $\mathbb{R}^{n}$ such that $Q\subseteq Q_{0}=[0,1]^{n}$ for all $Q\in{\mathcal{S}}$. We define the (pointwise) hypersingular sparse operator associated with $\mathcal{S}$ by
$\mathbb{A}_{\mathcal{S}}^{t}f(x):=\sum_{Q\in\mathcal{S}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t-1}}\langle|f|\rangle_{Q}=\sum_{Q\in\mathcal{S}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t}}\int_{Q}|f(y)|\,dy.$
Remark 5.3.
The normalization $Q_{0}=[0,1]^{n}$ is made only for convenience. In general, it suffices to assume that there exists a fixed dyadic cube $Q_{0}$ such that $Q\subseteq Q_{0}$ for all $Q\in{\mathcal{S}}$; by a translation and dilation one may then reduce to the above normalized situation.
This global containment assumption is natural in the present hypersingular regime. Indeed, as explained in Remark 1.5 (2), when $t>1$ the geometry of the ambient domain becomes relevant. On the other hand, from the viewpoint of sparse domination in harmonic analysis, the assumption is mild: most operators of interest (such as Calderón–Zygmund operators and Hilbert transforms along monomial curves) are local, or can be decomposed into a sum of localized pieces, and this yields the existence of such a global cube $Q_{0}$ (for each localized piece).
A natural question is the following.
Question 5.4.
Let $\eta\in(0,1)$, $t>1$ and $\mathcal{S}$ be an $\eta$-sparse family in $\mathbb{R}^{n}$ defined as above. For which pairs $1\leq p,q\leq\infty$ does $\mathbb{A}_{\mathcal{S}}^{t}$ extend to a bounded operator
$\mathbb{A}_{\mathcal{S}}^{t}:L^{p,1}(\mathbb{R}^{n})\to L^{q,\infty}(\mathbb{R}^{n})$
(or satisfy a stronger estimate at the pair $(p,q)$, e.g. a weak-type, or strong-type bound, and more generally a Lorentz-space bound)?
5.1. Graded family
It turns out that, in addition to the sparseness of $\mathcal{S}$}, there is another fundamental structural parameter that influences the behavior of the hypersingular sparse operator $\mathbb{A}_{{\mathcal{S}}}^{t}$.
The key observation is already contained in Example 1.8, which shows that one must control how the cube sizes in $\mathcal{S}$} change from one “layer” to the next. This motivates us to introduce the notion of a graded family of dyadic cubes.
We now turn to some details. Let ${\mathcal{G}}\subseteq{\mathcal{D}}$ be a collection of dyadic cubes. Again, we may assume $Q\subseteq[0,1]^{n}$ for all $Q\in{\mathcal{S}}$.
$\bullet$ First, let ${\mathcal{G}}^{(0)}$ denote the collection of all maximal dyadic cubes in $\mathcal{G}$}. For simplicity, we may assume that ${\mathcal{G}}^{(0)}=\{[0,1]^{n}\}$. Otherwise, we decompose $\mathcal{G}$} into finitely many such collections and treat each one separately (by translation and dilation). To this end, we define $\mathfrak{G}_{0}:=\ell(Q_{0})$, where $\ell(Q)$ denotes the sidelength of a dyadic cube $Q$.
$\bullet$ Next, let ${\mathcal{G}}^{(1)}$ be the subcollection of all maximal dyadic cubes in ${\mathcal{G}}\setminus{\mathcal{G}}^{(0)}$, and define $\mathfrak{G}_{1}:=\inf_{Q\in{\mathcal{G}}^{(1)}}\ell(Q)$. Iterating this procedure, we obtain a decomposition of $\mathcal{G}$} into layers $\{{\mathcal{G}}^{(j)}\}_{j\geq 0}$ together with the associated scales $\{\mathfrak{G}_{j}\}_{j\geq 0}$. Observe that for each $j\geq 1$, the dyadic cubes in ${\mathcal{G}}^{(j)}$ are mutually disjoint.
Definition 5.5.
Let ${\mathcal{G}}\subseteq{\mathcal{D}}$ be a collection of dyadic cubes in $\mathbb{R}^{n}$ such that ${\mathcal{G}}^{(0)}=\{[0,1]^{n}\}$, and write ${\mathcal{G}}=\bigcup_{j\geq 0}{\mathcal{G}}^{(j)}$ as above. We say that $\mathcal{G}$} is graded if
$K_{{\mathcal{G}}}:=\sup_{j\geq 0}\left(\log_{2}\frac{\mathfrak{G}_{j}}{\mathfrak{G}_{j+1}}\right)<\infty.$
We call $K_{{\mathcal{G}}}$ the degree of $\mathcal{G}$}, and we refer to ${\mathcal{G}}_{j}$ as the $j$-th layer of $\mathcal{G}$}.
Remark 5.6.
  1. (1)
    Here, we may assume that ${\mathcal{G}}_{1}\neq\emptyset$. Otherwise, $\mathbb{A}_{{\mathcal{G}}}^{t}f(x)=\mbox{1\hskip-4.25pt{1}}_{[0,1]^{n}}(x)\int_{[0,1]^{n}}f$, which is a rank-one operator and, in particular, maps $L^{1}(\mathbb{R}^{n})$ boundedly into $L^{\infty}(\mathbb{R}^{n})$.
  2. (2)
    Here are some examples of graded family. Let ${\mathcal{G}}^{(1)}$ consist of $0,1]^{n}$ together with all of its dyadic descendants; then ${\mathcal{G}}^{(1)}$ is graded (in fact, its degree is $1$). Another example is the collection [^5] In that setting, the role of $Q_{0}$ is replaced by $\mathbb{D}$}, and the sidelength $\ell(Q)$ in Definition 5.5 is replaced by the length of the boundary arc associated with the Carleson box. of dyadic Carleson boxes that appeared in our earlier analysis of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$. This family is also graded, again with degree $1$. Finally, we observe that $K_{{\mathcal{G}}}\geq 1$.
  3. (3)
    It is clear that a graded family need not be sparse, and vice versa.
  4. (4)
    It is not correct to replace the $\sup_{j\geq 0}$ in the definition of $K_{{\mathcal{G}}}$ by the quantity
    $K_{{\mathcal{G}}}^{\prime}:=\limsup_{k\to\infty}\left(\log_{2}\frac{\mathfrak{G}_{j}}{\mathfrak{G}_{j+1}}\right).$
    Indeed, $K_{{\mathcal{G}}}^{\prime}$ only controls the ratios $\mathfrak{G}_{j}/\mathfrak{G}_{j+1}$ for sufficiently large $j$ and, in particular, imposes no restriction on the initial scales (for instance, on $\mathfrak{G}_{0}/\mathfrak{G}_{1}$). Consequently, one loses uniform control on the gaps between the first few layers. Using the same idea as in Observation 1.8, one can construct a sequence of sparse families with the same sparseness and the same value of $K_{{\mathcal{G}}}^{\prime}$, but for which the associated hypersingular sparse operators still exhibit the “blow-up” phenomenon described in Example 1.8.
5.2. $L^{p}$ theory for hypersingular sparse operator $\mathbb{A}_{{\mathcal{S}}}^{t}$ induced by graded family
We have the following result.
Theorem 5.7.
Let $\eta\in(0,1)$, ${\mathcal{S}}\subseteq{\mathcal{D}}$ be a graded (contracting) $\eta$-sparse family in $0,1]^{n}$ with degree $K_{{\mathcal{S}}}$ as in Definition 5.5, $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$, and $\mathbb{A}_{{\mathcal{S}}}^{t}$ be the associated hypersingular sparse operator. Then the following statements hold.
  1. (1)
    (Off-critical line estimate) $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p}(\mathbb{R}^{n})$ to $L^{q}(\mathbb{R}^{n})$ when $(p,q)$ belongs to the off-critical line regime associated to $\mathbb{A}_{{\mathcal{S}}}^{t}$ given by
    (5.1) $\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}\right\}.$
  2. (2)
    (Critical line estimate) We consider two different cases.
    1. (a)
      $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p,\infty}(\mathbb{R}^{n})$ to $L^{q}(\mathbb{R}^{n})$ when $(p,q)$ belongs to
      (5.2) $\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}\quad\textrm{and}\quad q\neq 1\right\}.$
    2. (b)
      $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p,1}(\mathbb{R}^{n})$ to $L^{q,\infty}(\mathbb{R}^{n})$ when
      (5.3) $p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\qquad\textrm{and}\qquad q=1.$
Before we prove Theorem 5.7, we make some remarks.
Remark 5.8.
Theorem 5.7 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.7 yields an alternative proof of [12, Theorem 3]. Finally, we note that the endpoint weak-type estimate (5.3) fails in general (see, Proposition 4.2).
Remark 5.9.
We make a further remark on the sparseness parameter. In our set-up, the sparseness $\eta\in(0,1)$ should always be understood with respect to the underlying grid. For the dyadic grids (base $2$), it is convenient to encode $\eta$ by
$\kappa:=-\log_{2}(1-\eta),$
where the base $2$ reflects the dyadic structure.
This normalization is stable under changing the base of the grid. For example, consider the Carleson boxes associated with a triadic system on $\mathbb{T}$ (i.e. in (2.1) we replace dyadic arcs of length $2^{-j}$ by triadic arcs of length $3^{-j}$). Let $\mathcal{I}$ be a triadic arc, then the corresponding “upper” region is given by
$Q_{\mathcal{I}}^{\mathrm{up}}:=\left\{z\in{\mathbb{D}}:\ \frac{z}{|z|}\in\mathcal{I},\ 1-|\mathcal{I}|\leq|z|<1-\frac{|\mathcal{I}|}{3}\right\},$
and hence the collection of all Carleson boxes associated with a triadic system is sparse with sparseness $\eta=2/3$. Therefore, if we normalize the sparseness using the base of the grid, then
$\log_{3}\Bigl(1-\frac{2}{3}\Bigr)=1,$
which coincides with the dyadic normalization $\log_{2}(1-1/2)=1$.
Proof of Theorem 5.7.
For each $j\geq 0$, let ${\mathcal{S}}^{(j)}$ denote the $j$-th layer of $\mathcal{S}$} as in Definition 5.5, and without loss of generality, we may assume ${\mathcal{S}}^{(0)}=\{[0,1]^{n}\}$.
Observe that, by the sparseness assumption and contracting property of $\mathcal{S}$}, for any $j\geq 0$, one has
(5.4) $\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|\lesssim(1-\eta)^{j},$
We divide the proof into several steps.
Step I: Weak-type bounds at the point $\left(1/p,1/q\right)=\left(0,\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}\right)$. Our goal is to show the boundedness of
(5.5) $\mathbb{A}_{{\mathcal{S}}}^{t}:L^{\infty}(\mathbb{R}^{n})\to L^{\frac{-\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}(t-1)},\infty}(\mathbb{R}^{n}).$
Let $\alpha>0$ and $f\in L^{\infty}(\mathbb{R}^{n})$ with $\|f\|_{L^{\infty}(\mathbb{R}^{n})}=1$. Then, we have to estimate the size of the level set $E:=\left|\left\{x\in[0,1]^{n}:\left|\mathbb{A}_{{\mathcal{S}}}^{t}f(x)\right|>\alpha\right\}\right|$. Without loss of the generality, we may assume $\alpha$ is sufficiently large. We have the following observation.
(1). First, consider ${\mathcal{S}}^{(0)}=\{[0,1]^{n}\}$. Observe that
[…]
By the sparseness and contracting property of $\mathcal{S}$}, $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\bigcup_{Q\in{\mathcal{S}}^{(1)}}Q$, whose size is at most $1-\eta$.
(2). Next, we consider the next layer ${\mathcal{S}}^{(1)}$. Using the sparseness and contracting property of $\mathcal{S}$} again, we find that
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq 1+\frac{1}{|Q|^{t}}\int_{Q}|f(x)|dx\leq 1+|Q|^{1-t}\leq 1+2^{nK_{{\mathcal{S}}}(t-1)},\quad z\in[0,1]^{n}\backslash\left(\bigcup_{Q\in{\mathcal{S}}^{(2)}}Q\right)$
and $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\bigcup_{Q\in{\mathcal{S}}^{(2)}}Q$, whose size is at most $(1-\eta)^{2}$.
Iterating the above procedure, we see that for any $J\geq 0$, if
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq\sum_{\ell=0}^{J}2^{n\ell K_{{\mathcal{S}}}(t-1)}=C_{1}2^{nJK_{{\mathcal{S}}}(t-1)}-C_{2},$
where $C_{1},C_{2}>0$ are some absolute constants [^6] Here, we can take $C_{1}=\frac{2^{nK_{{\mathcal{S}}}(t-1)}}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$ and $C_{2}=\frac{1}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$. that only depend on $n,K_{{\mathcal{S}}}$ and $t$, then $z\in[0,1]^{n}\backslash\left(\bigcup_{Q\in{\mathcal{S}}^{(J+1)}}Q\right)$, whose size is at least $1-(1-\eta)^{J+1}$. Therefore, for any $\alpha>0$ sufficiently large, if $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)>C_{1}\alpha-C_{2}$, then
$z\in\bigcup_{Q\in{\mathcal{S}}^{(\widetilde{J}+1)}}Q$
for $\widetilde{J}>\frac{\log_{2}\alpha}{nK_{\mathcal{S}}(t-1)}$, which implies
$|E|$ $=\left|\left\{x\in[0,1]^{n}:\left|\mathbb{A}_{{\mathcal{S}}}^{t}f(x)\right|>C_{1}\alpha-C_{2}\right\}\right|$
$\lesssim(1-\eta)^{\frac{\log_{2}\alpha}{nK_{\mathcal{S}}(t-1)}}=\alpha^{\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}(t-1)}}.$
Thus, this gives
$(C_{1}\alpha-C_{2})|E|^{\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}}\lesssim\alpha\cdot\left(\alpha^{\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}(t-1)}}\right)^{\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)}}\simeq 1,$
which concludes the desired weak-type bound (5.5).
Step II: Strong-type bounds within the off-critical line regime (5.1): off-critical line estimates. Let $(p,q)$ be a pair satisfying
(5.6) $\frac{1}{q}-\frac{1}{p}>\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)},$
which, in particular, gives $p>q$. Without loss of generality, we may assume that $p<\infty$, as the case $p=\infty$ is analogous and we would like to leave the details to the interested reader.
Our goal in the second step is to show that $\mathbb{A}_{{\mathcal{S}}}^{t}f$ extends a bounded operator from $L^{p}(\mathbb{R}^{n})$ to $L^{q}(\mathbb{R}^{n})$. Decompose
$\mathbb{A}_{{\mathcal{S}}}^{t}=\sum_{j\geq 0}\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t},$
where
$\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t}f(x):=\sum_{Q\in{\mathcal{S}}^{(j)}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t}}\int_{Q}|f(y)|\,dy$
It suffices to show that there exists some absolute constant $C_{3}>0$, such that for each $j\geq 1$,
(5.7) $\left\|\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t}\right\|_{L^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})}\lesssim 2^{-C_{3}j}.$
Indeed, since the dyadic cubes in ${\mathcal{S}}^{(j)}$ are mutually disjoint, we have
$\left\|\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t}f\right\|^{q}_{L^{q}(\mathbb{R}^{n})}$ $=\int_{[0,1]^{n}}\left|\sum_{Q\in{\mathcal{S}}^{(j)}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t}}\int_{Q}|f(y)|\,dy\right|^{q}dx$
$=\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|^{1-tq}\left(\int_{Q}|f(y)|dy\right)^{q}$
$=\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|^{1-tq}|Q|^{q}\left(\frac{1}{|Q|}\int_{Q}|f(y)|dy\right)^{q}$
$\leq\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|^{1-tq}|Q|^{q}\left(\frac{1}{|Q|}\int_{Q}|f(y)|^{p}dy\right)^{\frac{q}{p}}$
(5.8) $=\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|^{q(1-t)}\cdot|Q|^{\frac{p-q}{p}}\left(\int_{Q}|f(y)|^{p}dy\right)^{\frac{q}{p}},$
where $p^{\prime}$ is the conjugate of $p$ satisfying $1/p+1/p^{\prime}=1$. Since $\mathcal{S}$} is a graded family with degree $K_{{\mathcal{S}}}$, we have for each $Q\in{\mathcal{S}}^{(j)}$, $|Q|\geq 2^{-jK_{{\mathcal{S}}}n}$. Therefore, using the assumption that $t>1$ and (5.4), we derive that
RHS of (5.2) $\leq 2^{jK_{{\mathcal{S}}}nq(t-1)}\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|^{\frac{p-q}{p}}\left(\int_{Q}|f(y)|^{p}dy\right)^{\frac{q}{p}}$
$\leq 2^{jK_{{\mathcal{S}}}nq(t-1)}\left(\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|\right)^{\frac{p-q}{p}}\left(\sum_{Q\in{\mathcal{S}}^{(j)}}\int_{Q}|f(y)|^{p}dy\right)^{\frac{q}{p}}$
$\leq 2^{jK_{{\mathcal{S}}}nq(t-1)}\cdot(1-\eta)^{\frac{(p-q)j}{p}}\left\|f\right\|_{L^{p}(\mathbb{R}^{n})}^{q}$
$=2^{\,j\left(K_{{\mathcal{S}}}nq(t-1)+\frac{p-q}{p}\log_{2}(1-\eta)\right)}\left\|f\right\|_{L^{p}(\mathbb{R}^{n})}^{q},$
which gives
(5.9) $\left\|\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t}\right\|_{L^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})}\lesssim 2^{j\left(K_{{\mathcal{S}}}n(t-1)+\frac{p-q}{pq}\log_{2}(1-\eta)\right)}.$
Note that by (5.6), we have $K_{{\mathcal{S}}}n(t-1)+\frac{p-q}{pq}\log_{2}(1-\eta)<0$. Hence (5.7) holds, which completes Step II.
Step III: Restricted weak-type estimate at $\left(1/p,1/q\right)=\left(\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},1\right)$: critical line estimate, part (b). In this step, we use Bourgain’s interpolation trick. On one hand side, letting $p=\infty$ and $q=1$ in (5.9), we see that
$\left\|\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t}\right\|_{L^{\infty}(\mathbb{R}^{n})\to L^{1}(\mathbb{R}^{n})}\lesssim 2^{j\left(K_{{\mathcal{S}}}n(t-1)+\log_{2}(1-\eta)\right)}.$
Observe that $nK_{{\mathcal{S}}}(t-1)+\log_{2}(1-\eta)<0$, which follows from the assumption $t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$. On the other hand, we have
$\left\|\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t}f\right\|_{L^{1}(\mathbb{R}^{n})}$ $=\int_{\mathbb{R}^{n}}\left|\sum_{Q\in{\mathcal{S}}^{(j)}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t}}\int_{Q}|f(y)|\,dy\right|dx$
$=\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|^{1-t}\int_{Q}|f(y)|dy$
$\leq 2^{jK_{{\mathcal{S}}}n(t-1)}\sum_{Q\in{\mathcal{S}}^{(j)}}\int_{Q}|f(y)|dy$
$\leq 2^{jK_{{\mathcal{S}}}n(t-1)}\left\|f\right\|_{L^{1}(\mathbb{R}^{n})},$
which gives
(5.10) $\left\|\mathbb{A}_{{\mathcal{S}}^{(j)}}^{t}\right\|_{L^{1}(\mathbb{R}^{n})\to L^{1}(\mathbb{R}^{n})}\lesssim 2^{jK_{{\mathcal{S}}}n(t-1)}.$
Applying now Lemma 4.3 with
$\beta_{1}=K_{S}n(t-1),\quad\beta_{2}=-nK_{{\mathcal{S}}}(t-1)-\log_{2}(1-\eta),\quad p_{1}=q_{1}=q_{2}=1,\quad\textrm{and}\quad p_{2}=\infty,$
which gives the boundedness of
(5.11) $\mathbb{A}_{{\mathcal{S}}}^{t}:L^{p,1}(\mathbb{R}^{n})\to L^{q,\infty}(\mathbb{R}^{n}),$
with
$\theta=\frac{\beta_{2}}{\beta_{1}+\beta_{2}}=\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},\qquad\frac{1}{q}=1,$
and
$\frac{1}{p}=\frac{\theta}{p_{1}}+\frac{1-\theta}{p_{2}}=\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)}.$
The proof of Step 3 is complete.
Step 4: Weak-type bounds on the critical line (5.2): critical line estimate, Part (a). The last part simply follows from an application of the off-diagonal Marcinkiewicz interpolation between (5.5) and (5.11).
The proof is complete. ∎
6. Some open problems
1. Upgrading restricted weak type to weak type at the endpoint. It would be interesting to revisit the endpoint mapping properties of $K_{2t}$ at
$p=\frac{1}{3-2t}\quad\text{and}\quad q=1.$
We ask whether the restricted weak-type bound at this exponent can be upgraded to a weak-type estimate. We expect that any such refinement, if true, would require genuinely complex-analytic input: one would need to exploit additional cancellation coming from the kernel $(1-z\overline{w})^{-2t}$, rather than working only with the positive kernel $|1-z\overline{w}|^{-2t}$. The latter corresponds to the hypersingular Berezin transform $K_{2t}^{+}$, which fails to be bounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$ (see, Proposition 4.2).
2. Further applications of Bourgain’s interpolation on critical lines. It would be desirable to further explore Bourgain’s interpolation method in establishing critical-line estimates for other hypersingular operators. For instance:
(a) Forelli–Rudin type operators. For $a,b,c\in\mathbb{R}$, define
$T_{a,b,c}f(z)=(1-|z|^{2})^{a}\int_{{\mathbb{B}}_{n}}\frac{(1-|w|^{2})^{b}}{(1-\langle z,w\rangle)^{c}}\,f(w)\,dV(w),$
and
$S_{a,b,c}f(z)=(1-|z|^{2})^{a}\int_{{\mathbb{B}}_{n}}\frac{(1-|w|^{2})^{b}}{|1-\langle z,w\rangle|^{c}}\,f(w)\,dV(w),$
where ${\mathbb{B}}_{n}$ is the unit ball in $\mathbb{C}^{n}$ and $dV$ is the normalized volume measure on ${\mathbb{B}}_{n}$. In [28], Zhao and Zhou characterized the strong-type $L^{p}_{\alpha}({\mathbb{B}}_{n})\to L^{q}_{\beta}({\mathbb{B}}_{n})$ bounds for these operators under various assumptions on the parameters $a,b,c,\alpha$, and $\beta$. Here, for $1\leq p<\infty$ and $1<\alpha<\infty$, the space $L^{p}_{\alpha}({\mathbb{B}}_{n})$ is defined with respect to the measure $dv_{\alpha}(z)=c_{\alpha}(1-|z|^{2})^{\alpha}\,dV(z)$, where $c_{\alpha}$ is chosen so that $v_{\alpha}({\mathbb{B}}_{n})=1$. A natural question is whether the method developed in the present paper can be applied to obtain critical-line estimates for Forelli–Rudin type operators in the hypersingular regime (see, e.g., [28, Theorem 1.1]).
(b) Other hypersingular averaging operators and forms. One may consider an $r$-th mean variant of the hypersingular sparse operator: for $t>1$ and $r\geq 1$, define
$\mathbb{A}_{{\mathcal{S}},r}^{t}f(z):=\sum_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(z)}{|Q|^{t-1}}\langle|f|\rangle_{Q,r}=\sum_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(z)}{|Q|^{t-1+\frac{1}{r}}}\Bigl(\int_{Q}|f|^{r}\Bigr)^{1/r},$
where $\langle|f|\rangle_{Q,r}:=\bigl(|Q|^{-1}\int_{Q}|f|^{r}\bigr)^{1/r}$. More generally, for $t>1$ and $1\leq r,s<\infty$, one may introduce the $(r,s)$-hypersingular sparse form
$\Lambda^{t}_{{\mathcal{S}};r,s}(f_{1},f_{2}):=\sum_{Q\in{\mathcal{S}}}|Q|^{2-t}\,\langle|f_{1}|\rangle_{Q,r}\,\langle|f_{2}|\rangle_{Q,s}.$
It would be interesting to understand how the parameters $t,r,s$, together with the geometry of $\mathcal{S}$}, affect the boundedness properties of these operators and forms.